# Hyperbolic regularization effects for degenerate elliptic equations

### v1

@@ 3.1. Setup and reduction
$\limsup_{j}\operatorname{diam}([\partial_{z}w_{j}](x_{j}))\leq\operatorname{diam}([\partial_{z}w_{\infty}](x_{0})).$
To pass from Theorem 3.1 to Theorem A, we can simply exploit Proposition 2.2 and write $K$ as the graph of a 1-Lipschitz map $h$. Hence it suffices to show Theorem 3.1. Lemma 2.2 tells us that we can, without loss of generality, assume that $h$ is defined in $\mathbb{C}$ and:
$\|h\|_{L^{\infty}({\mathbb{C}},{\mathbb{C}})}<+\infty,$ (3.2)
for some $\tilde{\sigma}$ fulfilling (2.9). Now the property of solving (1.8) for a Lipschitz map $w$ is equivalent to solving
$w_{\bar{z}}=h(w_{z})\text{ a.e. in }\mathbb{C}.$ (3.3)
Notice that, by Theorem 1.1, $w_{\infty}$ also solves (1.8)-(3.3). From now on we will consider $K$ to be the graph of $h$. Furthermore, by considering $W_{j}(x)\doteq w_{j}(x+x_{j})-w_{j}(x_{j})$ and $W_{\infty}(x)=w_{\infty}(x+x_{0})-w_{\infty}(x_{0})$ instead of $w_{j}$ and $w_{\infty}$, we can assume in addition that
$x_{j}=x_{0}=0\text{ and }w_{j}(0)=w_{\infty}(0)=0,\quad\forall j\in{\mathbb{N}}.$ (3.4)
By rescaling the domain, we can then still assume these maps are defined in $B_{1}$.
Finally, we reduce the proof of Theorem 3.1 to the following:
@@
Let $w_{j},w_{\infty}$ be as in Theorem 3.1. Let $C\subset\mathbb{R}^{2}$ be an open, bounded convex set with $\partial_{z}w_{\infty}](B_{r})\subset C$ for some $r>0$. Then, given $a\in(\overline{C})^{c}$, there exist $\delta$ and $J$ depending on $r$ and $a$ such that
$\partial_{z}w_{j}](B_{\delta})\subset B_{\frac{\operatorname{d}(a,C)}{2}}^{c}(a),\quad\forall j\geq J.$
Lemma $3.$
Under the simplifications (3.4), Theorem 3.2 implies Theorem 3.1.
Proof.
Let $C^{\prime}\Subset C^{\prime\prime}\subset\mathbb{R}^{2}$ be open convex sets containing $\partial_{z}w_{\infty}](0)$ with $\operatorname{diam}(C^{\prime\prime})\leq\operatorname{diam}([\partial_{z}w_{\infty}](0))+\alpha$, $\alpha>0$ to be chosen. Let $L=1+\sup_{j}\{\|Dw_{j}\|_{L^{\infty}}\}$. We can cover the compact set $(C^{\prime\prime})^{c}\cap\overline{B_{L}}$ with finitely many balls taken from the open cover given by the balls $B_{\frac{\operatorname{d}(a,C^{\prime})}{2}}(a)$, centered at $a\in(C^{\prime\prime})^{c}\cap\overline{B_{L}}$. Let $a_{i}$ be the centers of these balls, for $i=1,\dots,N$. By (1.12), we find $r>0$ such that $\partial_{z}w_{\infty}](B_{r})\subset C^{\prime}$. Hence, by Theorem 3.2 for each $i$ we find $\delta_{i}$ and $J_{i}$ for which
$\partial_{z}w_{j}](B_{\delta_{i}})\subset B_{\frac{\operatorname{d}(a_{i},C^{\prime})}{2}}^{c}(a_{i}),\quad\forall j\geq J_{i}.$
Let $\delta=\min_{i}\delta_{i}>0$ and $J=\max_{i}J_{i}<+\infty$. Then,
$\partial_{z}w_{j}](B_{\delta})\subset\bigcap_{i}B_{\frac{\operatorname{d}(a_{i},C^{\prime})}{2}}^{c}(a_{i})=\left(\bigcup B_{\frac{\operatorname{d}(a_{i},C^{\prime})}{2}}(a_{i})\right)^{c}\subset\left((C^{\prime\prime})^{c}\cap\overline{B_{L}}\right)^{c}=C^{\prime\prime}\cup(\overline{B_{L}})^{c},\quad\forall j\geq J.$
We also have, for all $j$, $\partial_{z}w_{j}](0)\subset[\partial_{z}w_{j}](B_{\delta})\subset\overline{B_{L}}$, and hence
$\operatorname{diam}([\partial_{z}w_{j}](0))\leq\operatorname{diam}(C^{\prime\prime})\leq\operatorname{diam}([\partial_{z}w_{\infty}](0))+\alpha,\quad\forall j\geq J.$
The arbitrariness of $\alpha$ concludes the proof. ∎
@@ 3.2.1. Topological degree
(3.5) is shown in [23, Theorem 5.31]. Through (3.5) and [23, Theorem 5.30], we also get (3.6).
@@
This is classical and can be found e.g. in [23, Theorem 2.3]. ∎
@@ 3.2.2. Quasiregular mappings
(1) can be found in [3, Theorem 5.1] (see also [1] for the precise expression of $p$). (2)-(3) can be found in [2, Corollary 5.5.2]. Let us show (4), following the same argument of [19, Lemma 4.3]. Observe that $\varphi$ is open by (2). Suppose by contradiction that there exist distinct $x_{1},x_{2}\in U\cap\varphi^{-1}(W)$ such that $\varphi(x_{1})=\varphi(x_{2})=y$. Then, taking a sufficiently small $r>0$ such that
@@
We state the stability result for quasiregular maps established in [32, Proposition 1]. In that reference, the statement is shown in $\mathbb{R}^{n}$, but we will only need it for planar maps. We derive from it Corollary 3.2.2, which is essentially contained in [32, Proposition 2].
@@
There are several ways to see that the assumption $M<+\infty$ in the previous proposition is automatically satisfied. One can use, for instance, Stoilow Factorization [2, Theorem 5.5.1] or more purely topological methods valid in all dimensions [51, Proposition 4.10(3)].
@@
Fix $\Lambda>0$ such that $|Du|+|A|\leq\Lambda$ for all $A\in\Gamma$ and a.e. $x\in B_{R}$. The finiteness of such $\Lambda$ and (3.12) imply that the map $u^{A}\doteq u-L_{A}$ is a nonconstant quasiregular mapping in $B_{R}$ for all $A\in\Gamma$. Thanks to Remark 3.2.2 we may apply Proposition 3.2.2 to $u^{A}$ at $x_{0}=0$. Recalling that $u^{A^{\prime}}(0)=0$ for all $A^{\prime}\in K$, this yields the existence of $r_{A}=r_{A}(u)\in(0,R)$ and $\delta_{A}=\delta_{A}(\varepsilon,\Lambda)>0$ such that
@@
Thus for all $A\in\Gamma$ we have $u^{A}(\partial B_{r})\cap\overline{B_{\delta r}}=\emptyset$, hence the degree $\deg(u^{A},B_{r},y)$ is well-defined for all $y\in\overline{B_{\delta r}}$. Moreover it is independent of $A\in\Gamma$ by path-connectedness of $\Gamma$ and homotopy invariance of the degree, see Proposition 3.2.1. ∎
@@ 3.2.3. Approximation of the problem
The solvability of this system can be found in [22, Proposition 2], where it is shown that if $H:B_{1}\times\mathbb{C}\to\mathbb{C}$ is a measurable function satisfying $H(z,0)=0$ for a.e. $z\in B_{1}$ and
@@
$H_{\varepsilon}(z,\xi)\doteq(1-\varepsilon)(h(\partial_{z}w+\xi)-h(\partial_{z}w))\text{ and }\sigma(z)=-\varepsilon h(\partial_{z}w).$
@@
By choosing $\delta$ small, we deduce the required uniform bound. Now we can use the boundedness of the Beurling transform [2, Theorem 4.5.3] to obtain interior $W^{1,p}$ bounds for all $p\geq 1$, uniformly in $\varepsilon>0$: indeed, for any test function $\eta\in C_{c}^{\infty}(B_{1})$, the derivative $(\eta w^{\varepsilon})_{\bar{z}}=\eta_{\bar{z}}w^{\varepsilon}+(1-\varepsilon)\eta h(w^{\varepsilon}_{z})$ is uniformly bounded in $L^{p}$ since $W^{1,2}(B_{1})\subset L^{p}(B_{1})$ and $h\in L^{\infty}({\mathbb{C}})$, and thus $D(\eta w_{\varepsilon})$ is uniformly bounded in $L^{p}$. We are only left to show the convergence of $w^{\varepsilon}$ to $w$. Using the boundary conditions,
@@
We write $K$ as the graph of $h$, as in (3.1)-(3.2)-(3.3). For simplicity, we let
$g:\mathbb{R}^{2}\to K\text{ be defined as }g(z)\doteq(z,h(z)),\text{ and }U\doteq g\left((\overline{C})^{c}\right),$
and we also let $\tau\doteq\frac{\operatorname{d}(a,C)}{2}$. We also recall (3.4).
Step 1: degree of $w_{j}-L_{A}$ for $A\in U$. We show the existence of $J\in{\mathbb{N}}$, $R\in(0,r)$ and $\alpha>0$ such that
$\quad\deg(w_{j}-L_{A},B_{R},y)=1\,,\quad\forall y\in\overline{B_{2\alpha}}\,,\;\forall A\in\overline{U}\,,\;\forall j\geq J\,.$ (3.15)
Recall from Remark 3.2.1 that this assertion implicitly entails $(w_{j}-L_{A})(\partial B_{R})\cap\overline{B_{2\alpha}}=\emptyset$. Thanks to the uniform convergence of $w_{j}$ to $w_{\infty}$ and Proposition 3.2.1, it suffices to show
$\quad\deg(w_{\infty}-L_{A},B_{R},y)=1\,,\quad\forall y\in\overline{B_{3\alpha}}\,,\;\forall A\in\overline{U}\,.$ (3.16)
From (1.9) and the facts that $Dw_{\infty}\in K$ a.e., $\overline{U}\subset K$ and $\overline{U}\cap[Dw_{\infty}](B_{r})=\emptyset$, we see that
$\det(Dw_{\infty}-A)\geq\varepsilon\quad\text{a.e. in }B_{r}\,,\quad\forall A\in\overline{U}\,$,
where $\varepsilon=\sigma(\operatorname{d}(\overline{U},[Dw_{\infty}](B_{r})))>0$. Noting that the set $\overline{U}\cap g(\overline{B_{M}})$ is compact and path-connected for any $M>0$ such that $\overline{C}\subset B_{M}$, we apply Corollary 3.2.2 and deduce the existence of $R=R(M)\in(0,r)$, $\delta=\delta(M)>0$ such that
@@
$\quad\deg(w_{\infty}-L_{A},B_{\rho},y)=1\quad\forall y\in\overline{B_{\rho}}\,,\;\forall\rho\in(0,r)\,,\;\forall A\in g(B_{M}^{c})\,$, (3.17)
apply this to $\rho=R$ and choose $3\alpha=\min\{\delta,R\}$, for $R,\delta$ given by this choice of $M$. To establish (3.17), let $M\geq 2\|h\|_{\infty}$, $z\in B_{M}^{c}$, so that $|z|\geq M\geq 2\|h\|_{\infty}$, and $A=g(z)=(z,h(z))$. Recalling (2.3), we have
@@
Assume moreover $M>16\operatorname{Lip}(w_{\infty})$. Then for $0<\rho<r$ we have $\|w_{\infty}\|_{L^{\infty}(B_{\rho})}<M\rho/16$ and we can invoke Proposition 3.2.1 to deduce
$\deg(w_{\infty}-L_{A},B_{\rho},y)=1\,,\quad\forall y\in\overline{B_{M\rho/16}}\,,\;\forall\rho\in(0,r)\,$,
which gives (3.17) if we impose in addition $M\geq 16$.
Step 2: Jordan curves and the degree of the approximating maps. For $t\in[0,1]$ and $j\in{\mathbb{N}}$ we define
$\gamma(t)\doteq\tau e^{2\pi it}+a,\quad\Gamma(t)\doteq g(\gamma(t))\,,\quad f_{j,t}=w_{j}-L_{\Gamma(t)}\,.$ (3.18)
Since $\Gamma(t)\in U$ for all $t\in[0,1]$, from (3.15) we infer
$\deg(f_{j,t},B_{R},y)=1\quad\text{ for all }y\in\overline{B_{2\alpha}}\,,\;j\geq J,\;t\in[0,1]\,.$ (3.19)
Now we fix $j\geq J$ and let $w_{j}^{\varepsilon}$ be the family parametrized by $\varepsilon\in(0,1)$ provided by Lemma 3.2.3 applied to $w=w_{j}$. Then we define the maps
$w_{j,\varepsilon,t}\doteq w_{j}^{\varepsilon}-L_{\Gamma_{\varepsilon}(t)},\qquad\Gamma_{\varepsilon}\doteq(\gamma,(1-\varepsilon)h(\gamma))\,$,
which are quasiregular for all $j\in{\mathbb{N}}$, $\varepsilon\in(0,1)$ and for all $t\in[0,1]$, compare Remark 3.2.2. For this fixed $j\geq J$, we see from Lemma 3.2.3 that $w_{j,\varepsilon,t}$ converges locally uniformly in $B_{1}$ to $f_{j,t}$ as $\varepsilon\to 0$, uniformly with respect to $t\in[0,1]$. Recall from Remark 3.2.1 that (3.19) contains the information $f_{j,t}(\partial B_{R})\cap\overline{B_{2\alpha}}=\emptyset$. Choosing $\varepsilon_{0}=\varepsilon_{0}(j)>0$ such that $\|w_{j,\varepsilon,t}-f_{j,t}\|_{L^{\infty}(B_{R})}<\alpha$ for all $\varepsilon\in(0,\varepsilon_{0})$, by Proposition 3.2.1 we infer
[…]
Further, recalling that $f_{j,t}(0)=0$ and $\operatorname{Lip}(f_{j,t})\leq C$, we may fix $\delta=\delta(C,R)>0$ such that
$B_{2\delta}\subset B_{R}\cap(f_{j,t})^{-1}({B_{\alpha/2}}),\quad\forall t\in[0,1],\forall j\in{\mathbb{N}}\,$,
hence, for a possibly smaller $\varepsilon_{0}=\varepsilon_{0}(j)$,
$B_{2\delta}\subset B_{R}\cap(w_{j,\varepsilon,t})^{-1}({B_{\alpha}}),\quad\text{ for all $t\in[0,1],0<\varepsilon\leq\varepsilon_{0}$.}$ (3.21)
We now claim that
$w_{j,\varepsilon,t}\text{ is injective on }B_{2\delta},\quad\text{ for all $t\in[0,1],0<\varepsilon\leq\varepsilon_{0}$.}$ (3.22)
@@
To see this, note first that $w_{j,\varepsilon,t}$ cannot be constant on $B_{R}$ by (3.20), so it has positive Jacobian a.e. in $B_{R}$ by Proposition 3.2.2(3). Recalling also from Lemma 3.2.3 that $w_{j,\varepsilon,t}\in W^{1,p}(B_{R})$ for all $p\geq 1$, we can apply (3.6) and deduce that, for $0<\varepsilon\leq\varepsilon_{0}$ and $t\in[0,1]$:
[…]
Through Proposition 3.2.2(4), we infer that $w_{j,\varepsilon,t}$ is injective on $w_{j,\varepsilon,t}^{-1}(B_{\alpha})\cap B_{R}$, hence on $B_{2\delta}$ by (3.21), thus proving (3.22).
Step 3: Separation and conclusion. We can finally conclude the argument. Fix for the moment $j\geq J$, where $J$ was found in Step $1$. By (3.22), we can write:
$w_{j,\varepsilon,t}(x+re_{1})\neq w_{j,\varepsilon,t}(x),\quad\forall\varepsilon\leq\varepsilon_{0},r<\delta,x\in B_{\delta}.$
The definitions yield
$w^{\varepsilon}_{j}(x+re_{1})-w^{\varepsilon}_{j}(x)\neq r\Gamma_{\varepsilon}(t)e_{1},\quad\forall\varepsilon\leq\varepsilon_{0},r<\delta,x\in B_{\delta}.$ (3.24)
Now we need to observe that $t\mapsto\Gamma_{\varepsilon}(t)e_{1}$ is a Jordan curve. Hence, it divides $\mathbb{R}^{2}$ into two connected components, the bounded one, $\omega_{\varepsilon}$, and the unbounded one, $\mathbb{R}^{2}\setminus\omega_{\varepsilon}$. From (3.24), we deduce that
$x\mapsto\frac{w_{j}^{\varepsilon}(x+re_{1})-w_{j}^{\varepsilon}(x)}{r}$
does not intersect this Jordan curve. Hence, for fixed $j\geq J,\varepsilon<\varepsilon_{0}$, for all $r<\delta$ and for every $x\in B_{\delta}$,

### v2

@@ 4. Elliptic inclusions and separation
$\limsup_{j}\operatorname{diam}([\partial_{z}w_{j}](x_{j}))\leq\operatorname{diam}([\partial_{z}w_{\infty}](x_{0})).$ (4.1)
As a consequence,
$\limsup_{j}\operatorname{diam}([Dw_{j}](x_{j}))\leq 2\operatorname{diam}([Dw_{\infty}](x_{0})).$ (4.2)
Remark 4.2.
Notice that we only need to show (4.1), as (4.2) follows from the fact that $w_{j}$ and $w$ solve (2.3) and $K$ is the graph of a 11-Lipschitz maps over the plane of conformal matrices, compare Proposition 3.2.
We reduce the proof of Theorem 4.1 to the following:
@@
Let $K\subset\mathbb{R}^{2\times 2}$ be compact and fulfill (2.4). Let $w_{0}\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ solve (2.3), and $C\subset\mathbb{R}^{2}$ an open, bounded, convex set such that $\partial_{z}w_{0}](B_{1})\subset C$. Then there exists $\delta=\delta(w_{0},K,C)>0$ such that
$\|w-w_{0}\|_{L^{\infty}(B_{1})}\leq\delta\quad\Longrightarrow\quad[\partial_{z}w](B_{\delta})\subset\overline{C}\,$,
for all $w\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ solving (2.3).
Proof of Theorem 4.1 from Theorem 4.3.
Let $\varepsilon>0$ and $C\subset\mathbb{R}^{2}$ be an open convex set containing $\partial_{z}w_{\infty}](x_{0})$ with $\operatorname{diam}(C)\leq\operatorname{diam}([\partial_{z}w_{\infty}](x_{0}))+\varepsilon$. By definition (2.7) of the essential range at $x_{0}$, there exists $r>0$ such that $B_{r}(x_{0})\subset B_{1}$ and $\partial_{z}w_{\infty}](B_{r}(x_{0}))\subset C$. Applying Theorem 4.3 to the rescaled map $w_{0}(x)=(w_{\infty}(x_{0}+rx)-w_{\infty}(x_{0}))/r$, and scaling back, we deduce the existence of $\delta\in(0,r)$ such that $\partial_{z}w](B_{\delta}(x_{0}))\subset\overline{C}$, for any $w$ solving (2.3) and such that $\|w-w_{\infty}\|_{L^{\infty}(B_{1})}\leq\delta$. Fixing $J\geq 0$ such that $\|w_{j}-w_{\infty}\|_{L^{\infty}(B_{1})}\leq\delta$ and $|x_{j}-x_{0}|\leq\delta/2$ for all $j\geq J$, we have therefore
$[\partial_{z}w_{j}](x_{j})\subset[\partial_{z}w_{j}](B_{\delta}(x_{0}))\subset\overline{C}\,$,
hence $\limsup_{j}\operatorname{diam}([\partial_{z}w_{j}](x_{j}))\leq\operatorname{diam}(\overline{C})=\operatorname{diam}(C)\leq\operatorname{diam}([\partial_{z}w_{\infty}](0))+\varepsilon$. Letting $\varepsilon\to 0$ concludes the proof of (4.1). As said in Remark 4, (4.2) follows from it directly. ∎
@@ 4.1.1. Topological degree
(4.3) is shown in [25, Theorem 5.31]. Through (4.3) and [25, Theorem 5.30], we also get (4.4). ∎
@@
This is classical and can be found e.g. in [25, Theorem 2.3]. ∎
@@ 4.1.2. Quasiregular mappings
(1) can be found in [3, Theorem 5.1] (see also [1] for the precise expression of $p$). (2)-(3) can be found in [2, Corollary 5.5.2]. Let us show (4), following the same argument of [21, Lemma 4.3]. Observe that $\varphi$ is open by (2). Suppose by contradiction that there exist distinct $x_{1},x_{2}\in U\cap\varphi^{-1}(W)$ such that $\varphi(x_{1})=\varphi(x_{2})=y$. Then, taking a sufficiently small $r>0$ such that
@@
We state the stability result for quasiregular maps established in [34, Proposition 1]. In that reference, the statement is shown in $\mathbb{R}^{n}$, but we will only need it for planar maps. We derive from it Corollary 4.1.2, which is essentially contained in [34, Proposition 2].
@@
There are several ways to see that the assumption $M<+\infty$ in the previous proposition is automatically satisfied. One can use, for instance, Stoilow Factorization [2, Theorem 5.5.1] or more purely topological methods valid in all dimensions [53, Proposition 4.10(3)].
@@
Fix $\Lambda>0$ such that $|Du|+|A|\leq\Lambda$ for all $A\in\Gamma$ and a.e. $x\in B_{R}$. The finiteness of such $\Lambda$ and (4.10) imply that the map $u^{A}\doteq u-L_{A}$ is a nonconstant quasiregular mapping in $B_{R}$ for all $A\in\Gamma$. Thanks to Remark 4.1.2 we may apply Proposition 4.1.2 to $u^{A}$ at $x_{0}=0$. Recalling that $u^{A^{\prime}}(0)=0$ for all $A^{\prime}\in\Gamma$, this yields the existence of $r_{A}=r_{A}(u)\in(0,R)$ and $\delta_{A}=\delta_{A}(\varepsilon,\Lambda)>0$ such that
@@
Thus for all $A\in\Gamma$ we have $u^{A}(\partial B_{r})\cap\overline{B_{\delta r}}=\emptyset$, hence the degree $\deg(u^{A},B_{r},y)$ is well-defined for all $y\in\overline{B_{\delta r}}$. Moreover, it is independent of $A\in\Gamma$ by path-connectedness of $\Gamma$ and homotopy invariance of the degree, see Proposition 4.1.1. ∎
@@ 4.1.3. Approximation of the problem
Thanks to Proposition 3.2 and Lemma 3.2 we may write $K\subset\hat{K}$, where $\hat{K}$ still fulfills (2.4) and is the graph of $h\colon{\mathbb{C}}\to{\mathbb{C}}$ satisfying
$\|h\|_{L^{\infty}({\mathbb{C}},{\mathbb{C}})}<+\infty,$ (4.12)
for some $\tilde{\sigma}$ fulfilling (3.9). For a Lipschitz map $w$ solving (2.3) we have therefore
$w_{\bar{z}}=h(w_{z})\text{ a.e. in }B_{1}.$ (4.13)
@@
The solvability of this system can be found in [24, Proposition 2], where it is shown that if $H:B_{1}\times\mathbb{C}\to\mathbb{C}$ is a measurable function satisfying $H(z,0)=0$ for a.e. $z\in B_{1}$ and
@@
$H_{\varepsilon}(z,\xi)\doteq(1-\varepsilon)(h(\partial_{z}w+\xi)-h(\partial_{z}w))\text{ and }\sigma(z)\doteq-\varepsilon h(\partial_{z}w).$
@@
By choosing $\delta$ small, we deduce uniform $L^{2}$ bounds on $Dw_{\varepsilon}$. Through Poincaré inequality, our boundary conditions and the fact that $w_{2}^{\varepsilon}$ has zero average, we get that $w_{\varepsilon}$ is equibounded in $L^{p}$ for any $p\in[1,+\infty)$, since $W^{1,2}(B_{1})\subset L^{p}(B_{1})$. Now we can use the boundedness of the Beurling transform [2, Theorem 4.5.3] to obtain interior $W^{1,p}$ bounds for all $p\geq 1$, uniformly in $\varepsilon>0$: indeed, for any $\eta\in C_{c}^{\infty}(B_{1})$, $(\eta w^{\varepsilon})_{\bar{z}}=\eta_{\bar{z}}w^{\varepsilon}+(1-\varepsilon)\eta h(w^{\varepsilon}_{z})$ is uniformly bounded in $L^{p}$ since $w^{\varepsilon}$ is and $h\in L^{\infty}({\mathbb{C}})$, and thus $D(\eta w_{\varepsilon})$ is uniformly bounded in $L^{p}$. We only need to show the convergence of $w^{\varepsilon}$ to $w$. Using the boundary conditions,
@@
Recall that $K\subset\hat{K}$, where $\hat{K}$ fulfills (2.4) and is the graph of $h$ satisfying (4.11)-(4.12)-(4.13). Without loss of generality we assume $w(0)=w_{0}(0)=0$. For brevity, recalling (3.4), let
$g:\mathbb{R}^{2}\to\hat{K}\text{ be defined as }g(z)\doteq(z,h(z)),\text{ and }\hat{K}_{C}\doteq g\left(\mathbb{R}^{2}\setminus C\right)\subset\hat{K}\,$.
Step 1: degree of $w-L_{A}$ for $A\in\hat{K}_{C}$. We show the existence of $\alpha>0$ and $0<r<1$ such that
$\deg(w-L_{A},B_{r},y)=1\,,\quad\forall y\in\overline{B_{2\alpha}}\,,\;\forall A\in\hat{K}_{C}\,$, (4.16)
provided $\|w-w_{0}\|_{L^{\infty}(B_{1})}\leq\alpha$. Recall that this assertion implicitly entails $(w-L_{A})(\partial B_{r})\cap\overline{B_{2\alpha}}=\emptyset$, see Remark 4.1.1. Thanks to Proposition 4.1.1, it suffices to show
$\quad\deg(w_{0}-L_{A},B_{r},y)=1\,,\quad\forall y\in\overline{B_{4\alpha}}\,,\;\forall A\in\hat{K}_{C}\,.$ (4.17)
From (2.4) and the facts that $Dw_{0}\in K$ a.e., $\hat{K}_{C}\subset\hat{K}$ and $\hat{K}_{C}\cap[Dw_{0}](B_{1})=\emptyset$, we see that
$\det(Dw_{0}-A)\geq\varepsilon\quad\text{a.e. in }B_{1}\,,\quad\forall A\in\hat{K}_{C}\,$,
where $\varepsilon=\sigma(\operatorname{d}(\hat{K}_{C},[Dw_{0}](B_{1})))>0$. Noting that the set $\hat{K}_{C}\cap g(\overline{B_{M}})$ is compact and path-connected for any $M>0$ such that $\overline{C}\subset B_{M}$, we apply Corollary 4.1.2 and deduce the existence of $r=r(M)\in(0,1)$, $\delta=\delta(M)>0$ such that
@@
$\quad\deg(w_{0}-L_{A},B_{\rho},y)=1\quad\forall y\in\overline{B_{\rho}}\,,\;\forall\rho\in(0,1)\,,\;\forall A\in g(B_{M}^{c})\,$, (4.18)
apply this to $\rho=r$ and choose $4\alpha=\min(\delta,r)$, for $r,\delta$ given by this choice of $M$. To establish (4.18), let $M\geq 2\|h\|_{\infty}$, $z\in B_{M}^{c}$, so that $|z|\geq M\geq 2\|h\|_{\infty}$, and $A=g(z)=(z,h(z))$. Recalling (3.3), we have
@@
Assume moreover $M>16\operatorname{Lip}(w_{0})$. Then for $0<\rho<1$ we have $\|w_{0}\|_{L^{\infty}(B_{\rho})}<M\rho/16$ and we can invoke Proposition 4.1.1 to deduce
$\deg(w_{0}-L_{A},B_{\rho},y)=1\,,\quad\forall y\in\overline{B_{M\rho/16}}\,,\;\forall\rho\in(0,1)\,$,
which gives (4.18) if we impose in addition $M\geq 16$. This proves Step 1.
From now on we fix $w\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ solving (2.3) and such that $\|w-w_{0}\|_{L^{\infty}(B_{1})}\leq\alpha$, so that (4.16) is satisfied. To conclude the proof of Theorem 4.3, the main idea is that the degree property (4.16) is not so far from implying that $w-L_{A}$ is injective in the preimage of $B_{\alpha}$, because $w-L_{A}$ is orientation preserving thanks to (2.4). This injectivity, for all $A\in\hat{K}_{C}$, would then imply strong constraints on the difference quotients of $w$ and eventually the desired conclusion that $Dw(x)\in\overline{C}$. In fact, that (4.16) implies injectivity would be true if $w-L_{A}$ were quasiregular, see Proposition 4.1.2. Hence we apply the strategy we just described, but at the level of the approximation $w^{\varepsilon}$ provided by Lemma 4.1.3.
Step 2: Approximation. We start by fixing $R>0$ such that $\overline{C}\subset B_{R}$ and $K\subset g(B_{R})$. In particular we have $\operatorname{Lip}(w)\leq M\doteq R+\|h\|_{\infty}$. Recalling also that $w(0)=0$ we infer, for $\eta\doteq\min(r/2,\alpha/(8M))>0$,
$B_{2\eta}\subset B_{r}\cap(w-L_{A})^{-1}(\overline{B_{\alpha/2}})\,,\quad\forall A\in\hat{K}_{C}\cap g(\overline{B_{R}})\,.$ (4.19)
Next we fix $A\in\hat{K}_{C}\cap g(\overline{B_{R}})$, and $a\in\overline{B_{R}}\setminus C$ such that $A=g(a)=(a,h(a))$. For $0<\varepsilon<1$ we let $w^{\varepsilon}$ be given by Lemma 4.1.3, and $A^{\varepsilon}=(a,(1-\varepsilon)h(a))$, so that the maps $w^{\varepsilon}-L_{A^{\varepsilon}}$ are quasiregular, see Remark 4.1.2. Since $w^{\varepsilon}\to w$ uniformly in $\overline{B_{r}}$ and $|A^{\varepsilon}-A|\leq\varepsilon\|h\|_{\infty}$, we may choose $\varepsilon_{0}>0$ such that
$\|(w^{\varepsilon}-L_{A^{\varepsilon}})-(w-L_{A})\|_{L^{\infty}(B_{r})}<\frac{\alpha}{2}\,,\quad\forall\varepsilon\in(0,\varepsilon_{0})\,.$
Note that $\varepsilon_{0}$ is independent of $A$. Next, we fix $\varepsilon\in(0,\varepsilon_{0})$ and combine this with (4.19) to obtain
$B_{2\eta}\subset B_{r}\cap(w^{\varepsilon}-L_{A^{\varepsilon}})^{-1}(\overline{B_{\alpha}})\,.$ (4.20)
Combining it also with (4.16), which, recall Remark 4.1.1, entails $(w-L_{A})(\partial B_{r})\cap\overline{B_{2\alpha}}=\emptyset$, gives
$\deg(w^{\varepsilon}-L_{A^{\varepsilon}},B_{r},y)=1\,,\quad\forall y\in\overline{{B_{\alpha}}}\,$, (4.21)
@@
thanks to Proposition 4.1.1. In particular, $w^{\varepsilon}-L_{A^{\varepsilon}}$ cannot be constant on $B_{r}$, hence it has positive Jacobian a.e. by Proposition 4.1.2(3). Recalling from Lemma 4.1.3 that $w^{\varepsilon}\in W^{1,p}(B_{r})$ for all $p\geq 1$, we can apply (4.4) and deduce
$N(w_{\varepsilon}-L_{A^{\varepsilon}},B_{r},y)=\deg(w_{\varepsilon}-L_{A^{\varepsilon}},B_{r},y)=1,\quad\text{for a.e. }y\in\overline{B_{\alpha}}\,.$
Through Proposition 4.1.2(4) and (4.20), we infer that
$w^{\varepsilon}-L_{A^{\varepsilon}}\text{ is injective on }B_{2\eta}\,,\quad\forall\varepsilon\in(0,\varepsilon_{0})\,,\;\forall A^{\varepsilon}\in g^{\varepsilon}(\overline{B_{R}}\setminus C)\,$, (4.22)
where $g^{\varepsilon}(z)=(z,(1-\varepsilon)h(z))$.
Step 3: Separation and conclusion. For any $x\in B_{\eta}$ and $0<t<\eta$ we deduce from (4.22) that $x$ and $x+te_{1}$ have different images through $w^{\varepsilon}-L_{A^{\varepsilon}}$, which can be rewritten as
$\frac{w^{\varepsilon}(x+te_{1})-w^{\varepsilon}(x)}{t}\neq A^{\varepsilon}e_{1}\,,\quad\forall\varepsilon\in(0,\varepsilon_{0})\,,\;\forall A^{\varepsilon}\in g^{\varepsilon}(\overline{B_{R}}\setminus C)\,$, (4.23)
For all $\varepsilon\in[0,1)$, the continuous map $\Phi^{\varepsilon}\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$ given by
$\Phi^{\varepsilon}(z)=g^{\varepsilon}(z)e_{1}=ze_{1}+(1-\varepsilon)h(z)e_{1}\,$,
is injective thanks to (4.11), and has closed image thanks to the boundedness of $h$ (4.12), hence $\Phi^{\varepsilon}$ is a homeomorphism of $\mathbb{R}^{2}$. Since $C$ and $\mathbb{R}^{2}\setminus\overline{B_{R}}$ are disjoint connected sets, we deduce that the complement of $\Phi^{\varepsilon}(\overline{B_{R}}\setminus C)$ has two connected components, $\Phi^{\varepsilon}(\mathbb{R}^{2}\setminus\overline{B_{R}})$ and $\Phi^{\varepsilon}(C)$. The image of the continuous map
$B_{\eta}\times(0,\eta)\ni(x,t)\mapsto\frac{w^{\varepsilon}(x+te)-w^{\varepsilon}(x)}{t}\,$,
lies therefore, by (4.23), in one of these two connected components: for any $\varepsilon\in(0,\varepsilon_{0})$ we have
either $\frac{w^{\varepsilon}(x+te_{1})-w^{\varepsilon}(x)}{t}\in\Phi^{\varepsilon}(C)\quad\forall(x,t)\in B_{\eta}\times(0,\eta)\,$,